# Neural Attention Search

### v2

@@ Abstract
We present Neural Attention Search (NAtS), a framework that automatically evaluates the importance of each token within a sequence and determines if the corresponding token can be dropped after several steps. This approach can efficiently reduce the KV cache sizes required by transformer-based models during inference and thus reduce inference costs. In this paper, we design a search space that contains three token types: (i) Global Tokens  will be preserved and queried by all the following tokens. (ii) Local Tokens  survive until the next global token appears. (iii) Sliding Window Tokens  have an impact on the inference of a fixed size of the next following tokens. Similar to the One-Shot Neural Architecture Search approach, this token-type information can be learned jointly with the architecture weights via a learnable attention mask. Experiments on both training a new transformer from scratch and fine-tuning existing large language models show that NAtS  can efficiently reduce the KV cache size required for the models while maintaining the models’ performance.
Machine Learning, ICML
@@ 1 Introduction
The ability to understand and infer from long-context information is crucial for many tasks such as long document summarization (Zhang et al., 2023a) and question answering (Kociský et al., 2018; Dasigi et al., 2021), code generation (Guo et al., 2023; Liu et al., 2024) or multi-round dialogues (Yi et al., 2024). Thanks to the ability to query the information from any position of the historical sequence, transformer-based large language models (Brown et al., 2020; Jiang et al., 2023; Anthropic, 2024; Grattafiori et al., 2024; Gemini & et al., 2024) extend their context length up to millions of tokens.
@@
Indeed, not all the tokens in a sequence are equally important. Many of them are redundant and do not contribute to the final output. Humans can recognize this information without pre-defined fixed rules and summarize or discard the context information into much smaller content. Transformers could also learn this ability implicitly: Many tokens in the attention map might only have very low weights (Zhang et al., 2023b) and only have little influence on the final predictions. However, as the transformer learns this information implicitly, we might not know how the important tokens would be distributed in the context. Selecting these tokens and recognizing the attention distributions might require extra human experts’ knowledge by either looking at the attention maps (Liu et al., 2023; Zhang et al., 2023b; Feng et al., 2024; Zhang et al., 2024) or applying specific fixed rules (Child et al., 2019; Beltagy et al., 2024; Xiao et al., 2023; Chen et al., 2024; Ge et al., 2024). Since this knowledge is already contained in the transformer models, we could also ask the model to evaluate the importance of each token and learn to predict the optimal type for the given input tokens automatically.
Unlike prior works that rely on human expertise or predefined rules to identify important tokens, we propose a novel approach to evaluate the importance of each token by assigning different roles to each of the tokens. E.g. some tokens will be preserved until the end, while other tokens might only survive for a short amount of time. These roles measure the importance of each token and determine if they would survive within the next few tokens. Rather than pre-define a set of fixed rules for each token, we ask the model to learn this information automatically. Finding the optimal role for each token is similar to the process of neural architecture search (Elsken et al., 2019a; White et al., 2023), where an optimal architecture is searched by the optimizer for a given task. Thereby, in this work, we jointly optimize the choice of each token and the model weights by constructing a learnable attention mask. Our approach is implicitly motivated by the one-shot neural architecture search approach (Pham et al., 2018; Liu et al., 2019; Dong & Yang, 2019), where the model parameter and architecture parameters are jointly optimized during the search process. However, our approach, Neural Attention Search (NAtS), searches for the optimal token roles jointly with the attention weights.
@@
    We propose Neural Attention Search (NAtS), a framework that automatically searches for the optimal roles for each token in the input sequence.
@@
    We propose different token roles in our search space that can be later combined to construct a learnable attention mask and then jointly optimized with the model weights in NAtS.
@@ 2.1 Attention maps for transformers
Transformers (Vaswani et al., 2017) computes the correlation between different tokens by mapping the input sequences into three variables, $Q$, $K$ and $V$. An attention map $A$ is first computed by pairing $Q$ and $K$ and then scaled and normalized with a softmax function. Finally, this attention map $A$ is multiplied by $V$. Additionally, a mask $M$ is attached to the Attention Map to guide the attention to focus only on certain tokens. This can be expressed as
$A$ $=\frac{QK^{T}}{\sqrt{\mathit{d}_{head}}}$ (1)
$O$ $=\mathrm{softmax}(A+M^{add})V$ (2)
@@
The additive attention mask $M^{add}_{i,j}\in\{-\inf,0\}$ controls if the transformer needs to construct the correlation between $Q_{i}$ and $K_{j}$. A widely used mask $M^{add}$ used in transformers are casual masks (Vaswani et al., 2017), where the upper parts of the mask are filled with $-\inf$ each token can only query the token information before them.
@@
Computing the attention map of a sequence with length $\mathit{L}$ requires a complexity of $\mathcal{O}(\mathit{L}^{2})$ for both time and memory. Techniques such as KV cache lowers this complexity to the $\mathcal{O}(\mathit{L}_{cache})$ during inference time with $\mathit{L}_{cache}$ being the length of the KV cache. However, as an expense, the cost of storing the KV cache gradually becomes unneglectable with the increasing size of large language models and the context length they could handle.
Many studies aim to reduce the KV cache size by sparsifying the attention map with pre-defined rules. For instance, Sparse Attention (Child et al., 2019) combines the local sliding windows and atrous attention to approximate the full attention maps; Longformer (Beltagy et al., 2024) further proposes to predefine a set of special tokens as global attention that have full access to any other tokens and can be quired by any other tokens; Streaming LLM (Xiao et al., 2023) shows that by preserving the initial tokens and the most recent tokens during inference time could recover the performance of window attention. DuoAttention (Xiao et al., 2024) further illustrates that only a small fraction of the attention heads are critical for processing long context information and therefore, only those retrieval heads should be preserved with the full attention. LongCoder (Guo et al., 2023) pre-defines a set of bridge tokens to summarize the information in between and memory tokens that will not be dropped afterwards. SepLLM (Chen et al., 2024) shows that the information contained in a segment can be compressed into a set of special tokens.
Alternatively, past attention maps can decide which tokens to evict or preserve. H2O (Zhang et al., 2023b) and Scissorhand (Liu et al., 2023) both identify the importance of each token based on their contributions to the attention maps and only preserve the most important tokens. Other approaches such as FastGen (Ge et al., 2024), SnapKV (Li et al., 2024), PyramidKV (Cai et al., 2024), and AadaKV (Feng et al., 2024) all applied pre-defined fixed rules to identify the important KV values. Additionally, instead of simply removing the unimportant tokens, we can also merge them into the existing tokens. CAM (Zhang et al., 2024) achieves this by merging the to-be-evicted tokens into the remaining tokens.
In contrast to the fine-tuning-free approaches above, alternatively, we can also fine-tune the target model to achieve the desired sparsity.  Kim et al. (2022) ask the model to learn to drop the tokens with lower attention scores. Anagnostidis et al. (2023) shows that this sparsity can also be learned through sparse sigmoid functions. SPARSEK (Lou et al., 2024) only selects a constant number of KV Paris for each query by introducing a differentiable SparseK operator. Landmark Attention (Mohtashami & Jaggi, 2023) insert landmark tokens after certain time steps and apply these landmark tokens to summarize the information of the tokens before that landmark token and recover. Furthermore, Dynamic Token Pooling (Nawrot et al., 2023) segments the input sequences as pooling layers to reduce the overall computational costs, while Dynamic Memory Compression (Nawrot et al., 2024) tries to accumulate the tokens within the same subsequence into one single token.
@@ 2.2 Neural Architecture Search
Designing a neural architecture for a specific task might require a lot of trial and error. Neural Architecture Search (NAS) automates this process by searching in a pre-defined search space (Elsken et al., 2019b). Previous work on NAS mainly focused on searching within a discrete search space by sampling a new architecture from the search space, evaluating its performance, and updating the optimizers (Zoph & Le, 2017; Zoph et al., 2018). However, training every network from scratch requires lots of GPU hours. One-Shot NAS (Pham et al., 2018) approaches instead share the same weights of the operators w.r.t. all the architectures in the search space. This allows to jointly optimize the architectures and weights. DARTS (Liu et al., 2019) and GDAS (Dong & Yang, 2019) further relax the discrete search space into continuous values to optimize the architecture parameters with gradient descent. The One-Shot NAS approach allows the optimizers to efficiently search for the optimal architecture within a huge search space. Similarly, NAtS  has multiple options for each token as the search space and is able to search for the optimal token types jointly with the model weights. However, unlike One-Shot NAS approaches that consider the optimization problem as a bilevel-optimization problem and optimize the model weights and architecture weights alternatively, we optimize the token state information and model weights within one forward and backward pass. This is similar to a mixture-of-expert (MOE) model (Shazeer et al., 2017; Fedus et al., 2023). However, instead of distributing data across all experts uniformly, we only select one expert for each token and assign all data to that expert.
@@ 3.1 Search Space Design
Not all the tokens in a sequence are equally important. Some of the tokens might contain important information, and they should be preserved that will be further queried by the following tokens. Some tokens might only contribute to the prediction for the next few tokens. Just like a paragraph is composed of multiple sentences, a sequence can be divided into multiple sub-sequences containing different information; some tokens might only be required within these sub-sequences. Hence, we design a search space for each token’s role within the sequence and ask the model to automatically learn the optimal role for each token in the sequence.
@@
We first define Global Token as tokens containing important information that need to be preserved for the following predictions. Liu et al. (2023); Zhang et al. (2023b) showed that only a small fraction of the tokens contributes to most of the attention scores for self-attention computation. These tokens need to be preserved for models to recall the global information that helps to predict the next token. In vanilla transformer models, all the tokens are Global Tokens.
@@
Global Tokens will not be evicted during inference time. Therefore, we should maintain as few Global Tokens as possible to ensure inference efficiency. Each Global Token should not only preserve the information within that position. Ideally, it should also be able to summarize the information from previous sequences (Guo et al., 2023; Chen et al., 2024). Hence, we split the entire sequence with the Global Tokens into multiple sub-sequences, with each sub-sequence ending with one Global Token. Each Global Token only needs to summarize the information from its sub-sequences and the previous Global Tokens.
@@
Local Token only survives until the next Global Token appears. Therefore, models will have the full attention within each sub-sequence to summarize the local sub-sequence information into the Global Token located at the end of the sub-sequence while being sparse within the input sequence.
@@
Only the Global Tokens and Local Tokens might control the sparsity at a low granularity level. E.g., assuming that one input sequence is highly localized, each token only has a high correlation with itself or a few neighboring tokens. In this case, they are all similar and are assigned with the same token type. However, none of the Global Token and Local Token could sparsify this attention map efficiently: if all the tokens are classified as Local Tokens, then the input sequence will only be considered as one single subsequence, and all the Local Tokens will be equivalent to the Global Tokens.
@@
Hence, we introduce Sliding Window Token. Sliding Window Tokens will only be preserved for the next W time steps and were previously considered as one of the most popular sparse attention approaches (Child et al., 2019; Xiao et al., 2023; Zhang et al., 2023b; Ge et al., 2024).
@@
In contrast to other causal attention maps, Figure LABEL:fig:LearnedAttention illustrates an exemplary attention mask constructed by the choices of different token types. In this case, we define the sliding window size as 4. Token 1, 4, 10 act as Global Tokens; Tokens 2, 6, 7, 8, 11 are Local Tokens; Token 3, 5, 9, 11 are Sliding Window Tokens. The Global Tokens splits the entire sequence into three subsequences that end at 4, 10, and the last index accordingly. Hence, Token 2 will only be quired by Token 3 and 4. This rule applies the same for Token 6, 7, and 8, where they only interact with the tokens within the same subsequence. For the sliding window tokens, only the next three tokens could query their information, regardless of whether these tokens belong to the same sub-sequence. Only 6 out of 12 tokens are involved during inference time to make the next token prediction.
@@ 3.2 Searching for the Optimal Token Types
Searching for the optimal token types within a sequence is similar to searching for the optimal architectures for a given task in neural architecture search (Elsken et al., 2019b; White et al., 2023). Following GDAS (Dong & Yang, 2019), we apply the Gumbel-Softmax trick (Jang et al., 2017; Maddison et al., 2017) to sample from the discrete search space. The Gumbel-Softmax trick allows the gradient to be backpropagated through the discrete choice of the token types.
@@
We could now construct a learnable attention mask $M$ with a series of sampled token states. However, the additive mask in Eq. 1 will take $-\inf$ values, resulting in invalid gradient information. Hence, we use the token information to construct a multiplicative attention mask $M^{mul}\in\{0,1\}$ [^1] For the sake of brevity, we will use $M$ instead of $M^{mul}$ in the following part of this paper.:
$O=\frac{e^{A}\odot M^{mul}}{\sum_{j}e^{A_{.,j}}\odot M^{mul}_{.,j}}V$ (3)
The attention mask columns for Global Tokens and Sliding Window Tokens can be directly constructed since they will survive for a fixed amount of steps. However, the mask for Local Tokens $M^{L}_{i,j}$ is controlled by both the distribution from Local Tokens and Global Tokens as Local Tokens will survive until the next Global Token appears. In other words, to make $M^{L}_{i,j}(j>i)$ a valid value, no Global Token should appear between $i$ and $j$.
@@
$M^{SW}_{i,j}$ =$=\begin{cases}1&\text{if}\ j\leq i+\textit{W}\\ 0&\text{if}\ j>i+\textit{W}\end{cases}$ (5)
@@
where W is the sliding window size and $G_{n}$ a Global Token at Position $n$. We could then construct the attention masks based on the type of each token. After that, we mask out the upper triangular part of the mask to ensure its causality.
In practice, we first collect the index of the next global token $K_{i}:=\min(\{k|k>=i\land G_{k}=1\})$ and rewrite Eq. 6 as:
$M^{L}_{i,j}=\begin{cases}1&\text{if}\ j\leq K_{i}\\ 0&\text{if}\ j>K_{i}\end{cases}$ (7)
These rules are then integrated in FlashAttention (Dao et al., 2022; Dao, 2024) to avoid explicitly computing the attention masks during the forward pass. In addition to the transformer computation, we only need to collect the next Global Token indices $\mathbf{K}$ (with a complexity of $\mathcal{O}(N)$) and then mask out the attention map values with the masks defined above.
$\frac{\partial O}{\partial M}=\frac{\partial O}{\partial g}\frac{\partial g}{\partial M}\qquad\frac{\partial g}{\partial M}=e^{A}\qquad\frac{\partial g}{\partial A}=e^{A}\odot M$ (8)
Hence, the gradients for $M$ are the same as those for $A$ if $M$ is 1. For the case of $M=0$, this is the same as the values that $\partial O/\partial A$ is supposed to be if no learnable mask is applied to the attention map. Hence, we can jointly compute the gradient computation from both M and A within one backward pass. Further details can be found in the appendix.
$d\alpha_{i}=\sum_{j}dM_{i,j}\times M_{i,j}^{\alpha}~{}$ (9)
Where $\alpha\in\{G,L,SW\}$ is the discretized token type. Intuitively, this shows the model’s preference over short-range or long-range correlations: If $i$ is quite close to $j$, then all the $\alpha$ will receive the same gradient information. However, if the model wants to create a long-range correlation with $i\gg j$, only Global Tokens will receive the gradient information and update their weights with the gradient.
In cases where $G_{k}$ is 0, Equation 10 is the negative value of Equation 6. However, for the case where $G_{k}=1$, this is equivalent to the local mask values where $k$ is no longer set as a Global Token. This requires us to find the index of the next global token $K_{k+1}$ and the last global token $K_{k-l}$ where $l$ is the length of the local sequence that ends at $k$.
$d\alpha^{G-}_{i}$ =$=\begin{cases}\sum\limits_{\begin{subarray}{c}m\geq i\\ n<i\end{subarray}}M^{L}_{m,n}\times dM_{m,n}\ \ &\text{if}\ G_{i}=0,\\ \sum\limits_{\begin{subarray}{c}K_{i+1}\geq m>=i\\ i>n\geq K_{k-l}\end{subarray}}dM_{m,n}\ \ &\text{if}\ G_{i}=1,\end{cases}$ (11)
$d\alpha^{G}_{i}$ $=d\alpha^{G}_{i}-d\alpha^{G-}_{i}$ (12)
To control the attention map’s sparsity, we introduce a small regularization value $\lambda$ that is directly applied to the gradient for Global Token and Local Tokens to encourage more Sliding Window Tokens. We show how $\lambda$ values control the attention map sparsity in the appendix.

### v3

@@ Abstract
We present Neural Attention Search (NAtS), an end-to-end learnable sparse transformer that automatically evaluates the importance of each token within a sequence and determines if the corresponding token can be dropped after several steps. To this end, we design a search space that contains three token types: (i) Global Tokens will be preserved and queried by all the following tokens; (ii) Local Tokens survive until the next global token appears; and (iii) Sliding Window Tokens have an impact on the inference of a fixed size of the next following tokens. Similar to the One-Shot Neural Architecture Search approach, this token-type information can be learned jointly with the architecture weights via a learnable attention mask. Experiments on both training a new transformer from scratch and fine-tuning existing large language models show that NAtS can efficiently reduce the KV cache size and the inference costs for the models while maintaining the models’ performance.
@@ 1 Introduction
The ability to understand and infer from long-context information is crucial for many tasks such as long document summarization [91] and question answering [19, 46], code generation [33, 54] or multi-round dialogues [88]. Thanks to the ability to query the information from any position of the historical sequence, transformer-based large language models [10, 13, 30, 32, 40] extend their context length up to millions of tokens.
@@
Indeed, not all the tokens in a sequence are equally important [44]. Many of them are redundant and do not contribute to the final output. Humans can recognize this information without pre-defined fixed rules and summarize or discard the context information into much smaller content. Transformers could also learn this ability implicitly: Many tokens in the attention map might only have very low weights [93] and only have little influence on the final predictions. However, as the transformer learns this information implicitly, we might not know how the important tokens would be distributed in the context. Selecting these tokens and recognizing the attention distributions might require extra human experts’ knowledge by either looking at the attention maps [25, 56, 92, 93] or applying specific fixed rules [12, 15, 16, 29, 82]. Since this knowledge is already contained in the transformer models, we could also ask the model to evaluate the importance of each token and learn to predict the optimal type for the given input tokens automatically.
Unlike prior works that rely on human expertise or predefined rules to identify important tokens [14, 25, 26, 29, 52, 82, 83, 93], we propose a novel approach to evaluate the importance of each token by assigning different roles to each of the tokens. For example, some tokens will be preserved until the end, while other tokens might only survive for a short amount of time. These roles measure the importance of each token and determine if they would survive within the next few tokens. Rather than pre-define a set of fixed rules for each token, we ask the model to learn this information automatically. Finding the optimal role for each token is similar to the process of neural architecture search [21, 79], where an optimal architecture is searched by the optimizer for a given task. Thereby, in this work, we jointly optimize the choice of each token and the model weights by constructing a learnable attention mask. Our approach is implicitly motivated by the one-shot neural architecture search approach [20, 53, 66], where the model parameters and architecture parameters are jointly optimized during the search process. However, our approach, Neural Attention Search (NAtS), searches for the optimal token roles jointly with the attention weights.
@@
    We propose Neural Attention Search (NAtS), an end-to-end learnable sparse attention framework that automatically searches for the optimal roles for each token in the input sequence.
@@
    We introduce different token roles in our search space that can be later combined to construct a learnable attention mask and then jointly optimized with the model weights in NAtS.
@@ 2.1 Attention maps for transformers
Transformers [77] computes the correlation between different tokens by mapping the input sequences into three variables, $Q$, $K$ and $V$. An attention map $A$ is first computed by pairing $Q$ and $K$ and then scaled and normalized with a softmax function. Finally, this attention map $A$ is multiplied by $V$. Additionally, a mask $M$ is attached to the Attention Map to guide the attention to focus only on certain tokens. This can be expressed as
$A=\frac{QK^{T}}{\sqrt{\mathit{d}_{head}}},\hskip 28.45274ptO=\mathrm{softmax}(A+M^{add})V$ (1)
@@
The additive attention mask $M^{add}_{i,j}\in\{-\inf,0\}$ controls if the transformer needs to construct the correlation between $Q_{i}$ and $K_{j}$. A widely used mask $M^{add}$ used in transformers is a causal mask [77], where the upper parts of the mask are filled with $\inf$; each token can only query the token information before it.
@@
Computing the attention map of a sequence with length $\mathit{L}$ requires a complexity of $\mathcal{O}(\mathit{L}^{2})$ for both time and memory. Techniques such as KV cache lower this complexity to the $\mathcal{O}(\mathit{L}_{\text{cache}})$ during inference time, with $\mathit{L}_{\text{cache}}$ being the length of the KV cache. However, as an expense, the cost of storing the KV cache gradually becomes unnegligible with the increasing size of large language models and the context length they could handle. Therefore, lots of work has been proposed to reduce the storage and computation overheads during the inference time [50].
To reduce the computational costs of an LLM on long context information, LLMLingua [64] trains another smaller model to compress the input texts. Many other studies work on reducing the KV cache size by deciding which tokens to evict or preserve with the past attention maps. Recent works such as H2O [93], Scissorhand [56], FastGen [29], KeyFormer [6], SnapKV [52], PyramidKV [14], AadaKV [25], ChunkKV [55], and CriticalKV [26] all apply different rules to identify the important of each token and remove all the remaining tokens. Additionally, instead of simply removing the unimportant tokens, we can also merge them into the existing tokens [78, 92].
Other works do not aim at reducing the KV cache size, but only select a subset of KV caches to compute the attention outputs, including Quest [74], TokenButler [8], AttentionPredictor [85], and XAttention [84]. However, these works still need to maintain the entire KV cache in the GPUs. This still results in a huge GPU memory usage overhead. To alleviate this issue, some other works like InfLLM [81], HIP [49], and ShodwKV [73] offload the KV cache to CPUs and only load the KV caches that are relevant for the current token predictions to GPUs.
@@
2.2 Sparse Attention
Sparse attention works on computing only a fraction of the attention map to reduce the computation costs and KV cache sizes with either pre-defined fixed patterns, including Sparse Transformers [16], Longformer [12], StreamingLLM [82], and DuoAttention [83]. Alternatively, this can be adjusted by special embedding tokens such as Longcoder [33] and SepLLM [15]. Additionally, one can also determine the types of sparse attention with a set of proxy values, including MInference [41], FlexPrefill [48], SeerAttention [28], SpargeAttention [90], and XAttention [84].
In contrast to the fine-tuning-free approaches above, we can also fine-tune the target model to achieve the desired sparsity. This includes Adaptively Sparse Attention [9] and Dynamic Memory Compression [62]. However, both approaches rely on expensive cumulative productions during training time, making their approach less applicable to long-context scenarios. LLMLingua2 [64] trains another model to compress the target texts. Additionally,  Kim et al. [44] asks the model to learn to drop the tokens with lower attention scores. Landmark Attention [61] inserts landmark tokens after certain time steps and applies these landmark tokens to summarize the information of the tokens before that landmark token and recover. SPARSEK [58] only selects a constant number of KV pairs for each query by introducing a differentiable SparseK operator. Mixture-of-Depth [70] only passes a limited number of tokens to each transformer layer. Native Sparse Attention [89] and MoBA [59] all select a subset of the KV values to compute the attention masks and therefore reduce the computational overhead. However, these approaches still require preserving the entire KV cache values and only reducing computational costs during decoding stages.
MoA [27] proposes to search for the optimal sparse attention for different heads, which share similar ideas to  NAtS. However, MoA’s search space only contains streaming LLM.This head-level search space of MoA is much smaller than our token-level search space, which might restrict the model’s expressibility under even smaller budgets.
@@ 2.3 Neural Architecture Search
Designing a neural architecture for a specific task might require a lot of trial and error. Neural Architecture Search (NAS) automates this process by searching in a pre-defined search space [22]. Previous work on NAS mainly focused on searching within a discrete search space by sampling a new architecture from the search space, evaluating its performance, and updating the optimizers [95, 96]. One-Shot NAS [66] approaches instead share the same weights of the operators with respect to all the architectures in the search space. This allows to jointly optimize the architectures and weights. DARTS [53] and GDAS [20] further relax the discrete search space into continuous values to optimize the architecture parameters with gradient descent. The One-Shot NAS approach allows the optimizers to efficiently search for the optimal architecture within a huge search space. Similarly, NAtS  has multiple options for each token as the search space and is able to search for the optimal token types jointly with the model weights. However, unlike One-Shot NAS approaches that consider the optimization problem as a bilevel optimization problem and optimize the model weights and architecture weights alternatively, we optimize the token state information and model weights within one forward and backward pass. This is similar to a mixture-of-experts (MOE) model [24, 71]. However, instead of distributing data across all experts uniformly, we only select one expert for each token and assign all data to that expert.
@@ 3.1 Search Space Design
Not all the tokens in a sequence are equally important. Just like a paragraph is composed of multiple sentences, a sequence can be divided into multiple sub-sequences containing different information; some tokens might only be required within these sub-sequences. Hence, we design a search space for each token’s role within the sequence and ask the model to automatically learn the optimal role for each token in the sequence.
@@
We first define Global Tokens as tokens containing important information that need to be preserved for the following predictions. Liu et al. [56] and Zhang et al. [93] showed that only a small fraction of the tokens contribute to most of the attention scores for self-attention computation. These tokens need to be preserved for models to recall the global information that helps to predict the next token. In vanilla transformer models, all the tokens are Global Tokens.
@@
Global Tokens will not be evicted during inference time. Therefore, we should maintain as few Global Tokens as possible to ensure inference efficiency. Each Global Token should not only preserve the information within that position. Ideally, it should also be able to summarize the information from previous sequences [15, 33]. Hence, we split the entire sequence with the Global Tokens into multiple sub-sequences, with each sub-sequence ending with one Global Token. Each Global Token only needs to summarize the information from its sub-sequences and the previous Global Tokens.
@@
Local Tokens only survive until the next Global Token appears. Therefore, models will have the full attention within each sub-sequence to summarize the local sub-sequence information into the Global Token located at the end of the sub-sequence while being sparse within the input sequence.
@@
Only the Global Tokens and Local Tokens might control the sparsity at a low granularity level. E.g., assuming that one input sequence is highly localized, each token only has a high correlation with itself or a few neighboring tokens. In this case, they are all similar and are assigned with the same token type. However, none of the Global Token and Local Token could sparsify this attention map efficiently: if all the tokens are classified as Local Tokens, the input sequence will only be considered as one single subsequence, and all the Local Tokens will be equivalent to the Global Tokens.
@@
Hence, we introduce Sliding Window Token. Sliding Window Tokens will only be preserved for the next W time steps and were previously considered as one of the most popular sparse attention approaches [16, 29, 82, 93].
(a) A comparison between different casual attention maps. LABEL:fig:FullAtt: The full attention map, where each token is connected to the tokens before it. LABEL:fig:LocalAtt: the local attention with sliding windows 3, every token will only get access to the information of the 3 tokens ahead. LABEL:fig:LongFormer Longformer, besides the local attention, the first, 6th and 9th tokens are the pre-defined global tokens. LABEL:fig:LearnedAttention: NAtS dynamically learns the optimal role for each token and constructs a learnable mask based on the tokens’ roles.
@@
In contrast to other causal attention maps, Figure LABEL:fig:LearnedAttention illustrates an exemplary attention mask constructed by the choices of different token types. In this case, we define the sliding window size as 4. Token 1, 4, 10 act as Global Tokens; Tokens 2, 6, 7, 8, 11 are Local Tokens; Token 3, 5, 9, 11 are Sliding Window Tokens. The Global Tokens splits the entire sequence into three subsequences that end at 4, 10, and the last index accordingly. Hence, Token 2 will only be required by Token 3 and 4. This rule applies the same for Token 6, 7, and 8, where they only interact with the tokens within the same subsequence. For the sliding window tokens, only the next three tokens could query their information, regardless of whether these tokens belong to the same sub-sequence. Only 6 out of 12 tokens are involved during inference time to make the next token prediction.
Combining different token types could already cover most of the sparse attention variables. For instance, assigning Global Tokens to the first few tokens and Sliding Window Tokens to all the other tokens results in streaming LLMs [82] and MoA [27]. If we further set all the tokens in some heads as Global Tokens, we will get DuoAttention [81]. Assigning different tokens as Global Tokens results in the Vertical patterns [41] that appear in many Token eviction strategies [52, 56]. Since different layers or heads might have different numbers of global tokens, patterns like pyramidkv [14] and adakv [25] can be easily derived from our search space. For a transformer network with $\mathit{N}$ layers and $\mathit{H}$ attention heads that needs to process one sequence of length $\mathit{L}$, given the three token types defined above, there are $3^{\mathit{L}\times\mathit{H}\times\mathit{N}}$ possible configurations in our search space. It is prohibitive to check all the combinations in such a huge search space, with the increasing sequence length that a model receives. Here, we propose an end-to-end approach to search for the optimal token types together with the network weights jointly.
@@ 3.2 Searching for the Optimal Token Types
Searching for the optimal token types within a sequence is similar to searching for the optimal architectures for a given task in neural architecture search [22, 79]. Following GDAS [20], we apply the Gumbel-Softmax trick [39, 60] to sample from the discrete search space. The Gumbel-Softmax trick allows the gradient to be backpropagated through the discrete choice of the token types.
Specifically, we first use a linear layer (which we call Attention Score Layer) that maps each the input tensor for an attention layer $X\in\mathbb{R}^{\mathit{d}}$ to the likelihood for each option: $\alpha\in\mathbb{R}^{(\mathit{H}*\mathit{N_{\text{opts}}})}=\text{Linear}(\mathbf{x})$, where $\mathit{H}$ is the number of KV heads and $\mathit{N_{\text{opts}}}$ is the number of options in the search space. The type $\alpha$ for each token is then sampled from a Gumble-Softmax function based on the likelihood values.
@@
$O=\frac{e^{A}\odot M^{\text{mul}}}{\sum_{j}e^{A_{.,j}}\odot M^{\text{mul}}_{.,j}}V$ (2)
The attention mask columns for Global Tokens and Sliding Window Tokens can be directly constructed since they will survive for a fixed number of steps. However, the mask for Local Tokens $M^{L}_{i,j}$ is controlled by both the distribution from Local Tokens and Global Tokens as Local Tokens will survive until the next Global Token appears. In other words, to make $M^{L}_{i,j}(j>i)$ a valid value, no Global Token should appear between $i$ and $j$.
@@
$M^{SW}_{i,j}$ $=\begin{cases}1&\text{if}\ j\leq i+\textit{W}\\ 0&\text{if}\ j>i+\textit{W}\end{cases}$ (4)
@@
where W is the sliding window size and $G_{n}$ a Global Token at Position $n$. We then construct the attention masks based on the type of each token. After that, we mask out the upper triangular part of the mask to ensure its causality.
In practice, we first collect the index of the next global token $I^{G_{\text{next}}}_{i}:=\min(\{k|k>=i\land G_{k}=1\})$ and rewrite Eq. 5 as:
$M^{L}_{i,j}=\begin{cases}1&\text{if}\ j\leq I^{G_{\text{next}}}_{i}\\ 0&\text{if}\ j>I^{G_{\text{next}}}_{i}\end{cases}$ (6)
During the backward process, we collect the column-wise gradient values from our attention masks and apply these gradients to update the parameters from our Attention Score Layer. Different from other approaches that approximates the attention maps or attention outputs from an existing model [8, 27, 85],  NAtS can be learned directly from the target labels, and thus, can be applied for both pre-training a new sparse transformer from scratch or fine tunning an existing transformer. To control the attention map’s sparsity, we introduce a small regularization value $\lambda$ that is directly applied to the gradient for Global Token and Local Tokens to encourage more Sliding Window Tokens. We show how the backward gradients is computed and how the $\lambda$ values control the attention map sparsity in the appendix  C.
These rules are then integrated in FlashAttention [17, 18] to avoid explicitly computing the attention masks during the forward pass. In addition to the transformer computation, we only need to collect the next Global Token indices $\mathbf{I^{G_{\text{next}}}}$ (with a complexity of $\mathcal{O}(N)$) and then mask out the attention map values with the masks defined above. At the same time, we also skip all the computation blocks that do not contain any valid values (i.e., all the items in the mask of that block are 0) during the online attention process. Further details can be found in the appendix D